OpenStudy (anonymous):

simplify: y+6=(5/6)(x-4)

OpenStudy (anonymous):

x=(6y+56)/5

OpenStudy (anonymous):

y=

OpenStudy (anonymous):

equation of the line; y= (sorry, forgot to state that)

OpenStudy (anonymous):

so which one is it? y+6=(5/6)(x-4) or x=(6y+56)/5 or are both part of the question?

OpenStudy (anonymous):

oh, thats someone elses answer y= (5/6)x + (8/3)

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

lishfosho?

OpenStudy (anonymous):

yes! thank you so much!